OpenStudy (anonymous):

Solve this using special products Foil method: (4x-1)(x-25) pls help me write here the answer i admitted it im SLOW

OpenStudy (sriram):

=(4x-1)x - (4x-1)25 =4x^2-x-100x+25 =4x^2-101x+25

OpenStudy (anonymous):

FOIL and unfortunate acronym for "first, outer, inner, last" something all math teachers will spend time in purgatory for. you have to do 4 multiplications here, First \[4x\times x\] Outer \[4x\times -25\]Inner \[-1\times x\] Last \[-1\times -25\]

OpenStudy (anonymous):

foil is just a memory aid for expanding these products first, outside, inside, last - referring to the position of the terms in the brackets.

OpenStudy (anonymous):

and a bad one at that!

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

when you do these 4 multiplications you get \[4x^2-100x-x+25\] then combine like terms to get \[4x^2-101x+25\]

OpenStudy (anonymous):

umm whats the answer? =4x^2-101x+25 is this?

OpenStudy (anonymous):

yes that is the answer